OpenStudy (anonymous):

In triangle PQR, angle P is congruent to angle R, ray of QT bisects angle PQR, and angle R=61 degrees. Find angle PQT. Justify your answer. Write in a statement reason format. @Mathematics

OpenStudy (anonymous):



OpenStudy (anonymous):

29

OpenStudy (anonymous):

see QT bisects the Angle PQR so Angle PQT is half of Angle PQR = 58/2 = 29

OpenStudy (anonymous):

Angle P = Angle R = 61 degrees So, the triangle is isosceles. 61 + 61 + PQR = 180 122 + PQR = 180 PQR = 180 - 122 = 58 So, PQT = 58/2 = 29 degrees.

OpenStudy (anonymous):

So how does that go in the 2 columns?